# Analyzing PDFs like Binaries: Adversarially Robust PDF Malware Analysis via Intermediate Representation and Language Model

### v1

@@ 4.1. Field Definitions
  • Index. This field indicates the index of the current IR within the object. Its value is uniquely determined by the combination of the object’s number and version, calculated as $<number-version>$. For instance, all IRs within the object 3 0 obj have an Index of 3-0, and within 8 2 obj, all IRs have an Index of 8-2.
@@ 4.2. Conversion
Basic Conversion Paradigm (P1) For basic object types such as numeric, string, name, boolean, null, and ID object, as mentioned in §$\S$2.1, the VType is assigned as num, str, name, bool, and null, respectively, with the Value maintaining its original form. For an indirectly referenced object, its type is assigned as ref, and the value is the index value of the referenced object. Figure 3 illustrates a basic conversion paradigm, which serves as the foundational transformation. All subsequent complex conversions are built upon this paradigm. The conversion process varies depending on whether the object is a stream, dictionary, single-element array, or mixed-element array, as will be discussed in detail later.
@@ 5.1. Preprocessing and Tokenization
In §$\S$4, we fixed the length of PDFObj IR and have already standardized it, so we do not require additional special tokens for normalization. To adapt PDFObj IR for pre-training, we need to perform tokenization. We employ the following tokenization strategy to mitigate the Out-Of-Vocabulary (OOV) issues caused by values: For each IR, we retain the Attribute and VType, and connect them with an underscore to form a single word. For example, in Figure 3, 7-0 Type name DictExp would be represented as Type_name. Multiple IRs form an object, and multiple words form a sentence; therefore, we treat an object as a sentence. The contextual relationships between sentences are determined by the reference relationships between objects. We extract these relationships from the ref, ref_list, and mix_list (where mix_list may include reference types) to generate the context of the sentences.
@@ 5.2. Pre-trained Mode
Masked Language Model To enable BERT to comprehend the internal structure of PDFObj IR, we first employed the Masked Language Model (MLM) training task. This task randomly masks tokens in the PDFObj IR text, forcing BERT to predict the masked content through bidirectional contextual inference. And it enables the model to learn deep semantic relationships between key-value pairs within objects. We began by pre-training BERT-based PDFObj2Vec using the MLM, following masking strategies from previous studies (Devlin et al., 2019; Li et al., 2021). For the input IR sequences $Seq_{IR}=IR_{1},IR_{2},...,IR_{i}$, in which $IR_{i}$ denotes a token, we randomly select 15% of the tokens to be masked. Of these tokens to be masked, 80% are replaced with the [MASK] token, 10% are replaced with a random IR token, and the remaining 10% are left unchanged. Subsequently, BERT’s encoder learns to predict the masked tokens:
@@ 5.3. General Mode
We also integrated three general embedding schemes to obtain embeddings for PDFObj IR: the BERT Base (Devlin et al., 2019; Research, 2018), CodeT5 (Wang et al., 2021), and text-embedding-3 (OpenAI, 2024). These general embedding models have been trained on extensive and diverse corpora. We downloaded the BERT Base and CodeT5 models to compute the embeddings for PDFObj IR. We also integrated the most popular conversational model, ChatGPT’s embedding model, text-embedding-3, which is OpenAI’s latest third-generation embedding model. We used the API provided by OpenAI to obtain IR embeddings directly. For preprocessing and tokenization, we adopted the same methods used in the previous pre-trained models.
@@ 6.1. Experimental Setup
Dataset We used the contagio (Parkour, 2013) dataset as our baseline dataset, which was commonly used in PDF malware analysis studies (Chen et al., 2020; Tong et al., 2019; Smutz and Stavrou, 2012). It has a balanced distribution of benign and malicious samples, comprising 9k benign samples and 11k malicious samples. Additionally, we collected an extended dataset for testing that includes 21k malicious samples from the CIC-PDFMal2022 dataset (Issakhani et al., 2022; for Cybersecurity, 2022) and 24k benign samples gathered from the internet. These benign samples include a diverse array of PDF types, such as bills, test files, books, and interactive forms, obtained from sources like GitHub, the PDF Association (Association, 2020), and the gov PDF dataset (of Congress, 2019). We used MD5 checksums to ensure that there was no overlap between the baseline and extended datasets. And we also used timestamps to confirm that samples in the extended dataset were collected after those in the baseline dataset.
@@
Baseline PDF Parsers To compare the parsing completeness with Poir, we select six popular PDF parsing tools, namely pdfrw (pmaupin/pdfrw, 2017), Poppler (Poppler, 2025), pdfminer (pdfminer/pdfminer.six, 2025), MuPDF (MuPDF, 2025), borb (jorisschellekens/borb, 2024), and QPDF (qpdf/qpdf, 2024), as baseline parsers. Among these six baseline parsers, pdfrw, pdfminer, and borb are written purely in Python, while Poppler and QPDF are written in C++, and MuPDF is written in C.
@@
PDFObj2Vec Configurations We implemented two modes of PDFObj2Vec: pre-trained and general. For the first mode, we pre-trained PDFObj2Vec schemes based on Word2Vec, PV-DM, and BERT, setting the embedding dimension of each model to 512. The specific training hyperparameters for these three pre-trained models are provided in Appendix D. For the general mode, we applied three off-the-shelf general embedding models: BERT Base (Devlin et al., 2019; Research, 2018), CodeT5 (Wang et al., 2021), and text-embedding-3 (OpenAI, 2024). BERT Base is a general embedding model for natural language, CodeT5 is a pre-trained embedding model for programming languages based on the T5 architecture, and text-embedding-3 is OpenAI’s latest third-generation embedding model. For these three schemes, we used their default embedding dimensions of 768, 256, and 1536, respectively.
@@
PDF Malware Classifiers Configurations We reproduced two state-of-the-art PDF robust malware classifiers (Chen et al., 2020; Tong et al., 2019), which demonstrate excellent adversarial robustness. One is robustly trained classifier (RTC). Chen et al. (Chen et al., 2020) used symbolic interval analysis to robustly retrain a deep neural network based on the robustness attributes defined by the structural paths. The other is conserved features trained classifier (CFTC). Tong et al. (Tong et al., 2019) adversarially retrained a support vector machine classifier based on the conservative features defined by the structural paths. Both RTC and CFTC’s reproduction details and hypermeters are provided in Appendix F. We followed the training setup on the baseline contagio dataset, as used by RTC and CFTC. Our implementation of the GIN-based classifier was also trained on this baseline dataset with hyperparameter details provided in Appendix E. All experiments were conducted on a machine with an Intel Core i7-12700K, NVIDIA GeForce GTX 3090, and 64 GB RAM running Ubuntu 20.04.
@@ 6.4. Evaluation of PDF Malware Classification
Classification Experiment Setup Following the settings of prior research (Chen et al., 2020; Tong et al., 2019), we splited the baseline dataset into training and test sets with a 7:3 ratio. We embedded the ORG nodes of the samples using various modes and schemes of PDFObj2Vec to generate AORG. Subsequently, we trained the GIN classifiers on the baseline training set, then we test and evaluate these GIN classifiers on both the baseline test set and the extended dataset. The comprehensive experimental results are summarized in Table 4.
@@
RTC (Chen et al., 2020) 99.15 99.97 98.11 85.15 97.04 74.72
CFTC (Tong et al., 2019) 97.58 99.75 95.04 92.04 97.31 87.41
@@
General Embedding Modes We evaluated three widely-used, general embedding models: BERT Base, CodeT5, and text-embedding-3, on the PDFObj IR to assess the effectiveness of applying general embeddings directly to classification. The results are presented in the last three rows of Table 4. The experimental results show that among the models tested on PDFObj IR, text-embedding-3 achieves the best performance in the downstream task of PDF malware classification. This can be attributed to its pre-training on a large-scale corpus, which provides it with numerous parameters that enhance its capability. However, text-embedding-3 is computationally expensive and cannot be deployed locally. In contrast, both BERT Base and CodeT5 are more cost-effective and can be deployed locally. Additionally, since CodeT5 is specifically pre-trained for programming languages, and PDFObj IR represents program-like structures, CodeT5 outperforms the general BERT model in PDF malware classification tasks. In the pre-training mode of PDFObj2Vec, the BERT scheme, which is specifically pre-trained for PDFObj IR, outperforms both the general BERT Base and CodeT5 models in downstream tasks. Although it slightly lags behind text-embedding-3 in overall accuracy, it performs better against adversarial attacks, a point we will discuss in the next subsection. Furthermore, we also explored the performance of directly embedding the raw content of PDF objects using these general embedding models. We evaluated their classification performance under the same experimental setup, and we will discuss these results in our ablation study (see §$\S$6.6).
@@ 6.5. Adversarial Attack
Gradient-Based Attack We adapted gradient-based white-box attack method (GradArgmax) from graph classification tasks (Dai et al., 2018; Wan et al., 2021) for PDF classification based on AORG. GradArgmax attack uses classifier gradient information to generate adversarial samples, aiming to reduce the model’s confidence in classifying the target sample. This approach evaluates the robustness of the integrated PDFObj2Vec graph classifier under fine-tuned adversarial examples. In this scenario, we assume the attacker has full access to the model’s gradient information, with no perturbation distance constraints. The attack budget is set to 1000, consistent with other graph attack study (Wan et al., 2021). The goal is to identify which graph edges most influence the target classification by evaluating their effect on the loss function using greedy strategy. Specifically, for each candidate edge, its gradient value is computed: if negative, the edge is considered for removal; if positive, it is considered for addition. Since the number of nodes and edges in AORG varies, smaller graphs may require only slight perturbations, while larger graphs need more significant ones. To address this, we introduce the relative perturbation ratio (RPR), defined as the ratio of perturbations to the maximum number of edges in the AORG.
@@
Genetic Algorithm-Based Attack We implemented a genetic algorithm-based black-box attack (GeneticAlg) (Dai et al., 2018), where the attacker has no access to the model’s internal details, only the output confidence scores. GeneticAlg leverages a genetic algorithm to optimize the structure of the input graph, effectively bringing the sample closer to the model’s decision boundary. This approach is used to evaluate the defensive capabilities of our proposed graph classifier against adversarial samples generated through complex search strategies. And, GeneticAlg consists of five key components: population, fitness function, selection, crossover, and mutation. The attack involves edge flipping, node injection, and deletion. Each generation evolves based on the fitness of the mutated samples, guiding the direction for subsequent generations. Following the experimental setup in (Dai et al., 2018), we set the attacker query limit to 1000, the population size to 100, and the number of generations to 10. The results, shown in Figure9(b), indicate that Word2Vec-20k and PV-DM-20k performed the worst in terms of TRA, while the other models maintained TRA values above 0.98, with text-embedding-3 and BERT-65k performing the best. Overall, the results suggest that our proposed classifier demonstrates strong defenses against the GeneticAlg attack.
@@
Random Noise Attack We extended node classification attack methods from (Ma et al., 2020) to perturb node features for whole-graph classification. Specifically, we injected Gaussian noise into the node features of the AORG graph to evaluate the model’s resistance to random perturbations, thereby testing its robustness under noisy conditions. To select the attack nodes, we employed degree centrality, as it provides a reasonable baseline for graph-based attacks (Newman, 2018). High-degree nodes were prioritized because they play a crucial role in information propagation and are more likely to significantly impact the overall graph structure. Gaussian noise was then introduced to perturb the features of these nodes. In this black-box attack scenario, the attacker only has access to the model’s binary output (0 or 1) and is limited to a maximum of 1000 queries. The experimental results, shown in Figure 9(c), indicate that Word2Vec exhibited the lowest robustness under these conditions. In contrast, BERT-65k and BERT-20k showed the highest performance, with TRA values of 1 and 0.999, respectively. It suggests that BERT scheme are more resilient to adversarial noise, particularly with respect to node feature perturbations. The TRA of text-embedding-3 showed a significant decline under the random noise attack, compared to the GradArgmax and GeneticAlg attacks. This indicates that its embeddings are highly susceptible to noise interference.
@@
Reverse Mimicry Attack We used the reverse mimicry adversarial attack (Maiorca et al., 2013) for evaluation, which has been evaluated in previous state-of-the-art robust PDF malware classifiers (Chen et al., 2020; Tong et al., 2019). This black-box attack is independent of specific classifiers or features, making it suitable for evaluating our proposed graph classifier as well as the RTC and CFTC classifiers, which employ different models and features. Unlike other attacks, reverse mimicry directly manipulates the sample space by injecting malicious payloads into benign samples, thereby generating realizable adversarial examples designed to evade classification boundaries. This approach enables a rigorous evaluation of the model’s robustness against intentionally disguised adversarial samples.
@@
We followed the adversarial evaluation settings from (Chen et al., 2020; Tong et al., 2019), using 500 seed samples to create adversarial samples. A Cuckoo sandbox (Sandbox, 2018) was set up to test the adversarial examples. If they generated the expected network communication metrics, they were deemed legitimate. The experimental results, shown in Figure 10, reveal that with a 20k corpus size, only BERT-20k demonstrated some resistance to adversarial samples. However, with a 65k corpus size, all three PDFObj2Vec pre-training models successfully detected adversarial samples. BERT-65k achieved a TRA of 1, while the other models had TRA values below 0.2. These findings suggest that BERT-65k is particularly effective at capturing the semantic information of nodes. When malicious payloads are injected into benign samples, the corresponding malicious nodes are also present in the adversarial samples. BERT-65k accurately represents these embeddings and, through the node propagation and aggregation process in GIN, successfully identifies the adversarial samples. Notably, RTC and CFTC achieved a TRA of over 0.5 but at the cost of significantly reduced classification accuracy, as shown in Table 4. Specifically, CFTC achieved only 97.58% accuracy on the baseline dataset. And both CFTC and RTC exhibited significantly lower TNR compared to our proposed classifier. This indicates a higher FPR for CFTC and RTC, for instance, CFTC reached an FPR of 4.96% on the baseline dataset and an FPR of 12.59% on the extended dataset, substantially higher than that of BERT-65k.
@@
Results Analysis We evaluated our proposed classifier using four adversarial attack methods that target different space. Among these, the reverse mimicry attack operates in the sample space and simulates a realistic black-box scenario. In contrast, GradArgmax, GeneticAlg, and RandomNoise are feature-space attacks that, while less realistic, are widely used to stress-test classifiers under strong assumptions (Dai et al., 2018; Wan et al., 2021; Ma et al., 2020; Chen et al., 2020). These attacks help reveal how vulnerable a model may be when its decision boundaries are exploited, offering valuable insights into worst-case robustness.

### v2

@@ 4.1. Field Definitions
  • Index. This field indicates the index of the current IR within the object. Its value is uniquely determined by the combination of the object’s number and version, calculated as <numberversion><number-version>. For instance, all IRs within the object 3 0 obj have an Index of 3-0, and within 8 2 obj, all IRs have an Index of 8-2.
@@ 4.2. Conversion
Basic Conversion Paradigm (P1) For basic object types such as numeric, string, name, boolean, null, and ID object, as mentioned in $\S$2.1, the VType is assigned as num, str, name, bool, and null, respectively, with the Value maintaining its original form. For an indirectly referenced object, its type is assigned as ref, and the value is the index value of the referenced object. Figure 3 illustrates a basic conversion paradigm, which serves as the foundational transformation. All subsequent complex conversions are built upon this paradigm. The conversion process varies depending on whether the object is a stream, dictionary, single-element array, or mixed-element array, as will be discussed in detail later.
@@ 5.1. Preprocessing and Tokenization
In $\S$4, we fixed the length of PDFObj IR and have already standardized it, so we do not require additional special tokens for normalization. To adapt PDFObj IR for pre-training, we need to perform tokenization. We employ the following tokenization strategy to mitigate the Out-Of-Vocabulary (OOV) issues caused by values: For each IR, we retain the Attribute and VType, and connect them with an underscore to form a single word. For example, in Figure 3, 7-0 Type name DictExp would be represented as Type_name. Multiple IRs form an object, and multiple words form a sentence; therefore, we treat an object as a sentence. The contextual relationships between sentences are determined by the reference relationships between objects. We extract these relationships from the ref, ref_list, and mix_list (where mix_list may include reference types) to generate the context of the sentences.
@@ 5.2. Pre-trained Mode
Masked Language Model To enable BERT to comprehend the internal structure of PDFObj IR, we first employed the Masked Language Model (MLM) training task. This task randomly masks tokens in the PDFObj IR text, forcing BERT to predict the masked content through bidirectional contextual inference. And it enables the model to learn deep semantic relationships between key-value pairs within objects. We began by pre-training BERT-based PDFObj2Vec using the MLM, following masking strategies from previous studies (devlin2019bert; li2021palmtree). For the input IR sequences $Seq_{IR}=IR_{1},IR_{2},...,IR_{i}$, in which $IR_{i}$ denotes a token, we randomly select 15% of the tokens to be masked. Of these tokens to be masked, 80% are replaced with the [MASK] token, 10% are replaced with a random IR token, and the remaining 10% are left unchanged. Subsequently, BERT’s encoder learns to predict the masked tokens:
@@ 5.3. General Mode
We also integrated three general embedding schemes to obtain embeddings for PDFObj IR: the BERT Base (devlin2019bert; standard_bert), CodeT5 (wang-etal-2021-codet5), and text-embedding-3 (openai_embedding). These general embedding models have been trained on extensive and diverse corpora. We downloaded the BERT Base and CodeT5 models to compute the embeddings for PDFObj IR. We also integrated the most popular conversational model, ChatGPT’s embedding model, text-embedding-3, which is OpenAI’s latest third-generation embedding model. We used the API provided by OpenAI to obtain IR embeddings directly. For preprocessing and tokenization, we adopted the same methods used in the previous pre-trained models.
@@ 6.1. Experimental Setup
Dataset We used the contagio (contagio13_online) dataset as our baseline dataset, which was commonly used in PDF malware analysis studies (chen2020training; tong2019improving; smutz2012malicious). It has a balanced distribution of benign and malicious samples, comprising 9k benign samples and 11k malicious samples. Additionally, we collected an extended dataset for testing that includes 21k malicious samples from the CIC-PDFMal2022 dataset (issakhani2022pdf; CICdataset) and 24k benign samples gathered from the internet. These benign samples include a diverse array of PDF types, such as bills, test files, books, and interactive forms, obtained from sources like GitHub, the PDF Association (stressfulpdf), and the gov PDF dataset (dotgovpdf). We used MD5 checksums to ensure that there was no overlap between the baseline and extended datasets. And we also used timestamps to confirm that samples in the extended dataset were collected after those in the baseline dataset.
@@
Baseline PDF Parsers To compare the parsing completeness with Poir, we select six popular PDF parsing tools, namely pdfrw (pdfrw), Poppler (Poppler55), pdfminer (pdfminer), MuPDF (mupdf), borb (borb), and QPDF (QPDF), as baseline parsers. Among these six baseline parsers, pdfrw, pdfminer, and borb are written purely in Python, while Poppler and QPDF are written in C++, and MuPDF is written in C.
@@
PDFObj2Vec Configurations We implemented two modes of PDFObj2Vec: pre-trained and general. For the first mode, we pre-trained PDFObj2Vec schemes based on Word2Vec, PV-DM, and BERT, setting the embedding dimension of each model to 512. The specific training hyperparameters for these three pre-trained models are provided in Appendix D. For the general mode, we applied three off-the-shelf general embedding models: BERT Base (devlin2019bert; standard_bert), CodeT5 (wang-etal-2021-codet5), and text-embedding-3 (openai_embedding). BERT Base is a general embedding model for natural language, CodeT5 is a pre-trained embedding model for programming languages based on the T5 architecture, and text-embedding-3 is OpenAI’s latest third-generation embedding model. For these three schemes, we used their default embedding dimensions of 768, 256, and 1536, respectively.
@@
PDF Malware Classifiers Configurations We reproduced two state-of-the-art PDF robust malware classifiers (chen2020training; tong2019improving), which demonstrate excellent adversarial robustness. One is robustly trained classifier (RTC). Chen et al. (chen2020training) used symbolic interval analysis to robustly retrain a deep neural network based on the robustness attributes defined by the structural paths. The other is conserved features trained classifier (CFTC). Tong et al. (tong2019improving) adversarially retrained a support vector machine classifier based on the conservative features defined by the structural paths. Both RTC and CFTC’s reproduction details and hypermeters are provided in Appendix F. We followed the training setup on the baseline contagio dataset, as used by RTC and CFTC. Our implementation of the GIN-based classifier was also trained on this baseline dataset with hyperparameter details provided in Appendix E. All experiments were conducted on a machine with an Intel Core i7-12700K, NVIDIA GeForce GTX 3090, and 64 GB RAM running Ubuntu 20.04.
@@ 6.4. Evaluation of PDF Malware Classification
Classification Experiment Setup Following the settings of prior research (chen2020training; tong2019improving), we splited the baseline dataset into training and test sets with a 7:3 ratio. We embedded the ORG nodes of the samples using various modes and schemes of PDFObj2Vec to generate AORG. Subsequently, we trained the GIN classifiers on the baseline training set, then we test and evaluate these GIN classifiers on both the baseline test set and the extended dataset. The comprehensive experimental results are summarized in Table 4.
@@
RTC (chen2020training) 99.15 99.97 98.11 85.15 97.04 74.72
CFTC (tong2019improving) 97.58 99.75 95.04 92.04 97.31 87.41
@@
General Embedding Modes We evaluated three widely-used, general embedding models: BERT Base, CodeT5, and text-embedding-3, on the PDFObj IR to assess the effectiveness of applying general embeddings directly to classification. The results are presented in the last three rows of Table 4. The experimental results show that among the models tested on PDFObj IR, text-embedding-3 achieves the best performance in the downstream task of PDF malware classification. This can be attributed to its pre-training on a large-scale corpus, which provides it with numerous parameters that enhance its capability. However, text-embedding-3 is computationally expensive and cannot be deployed locally. In contrast, both BERT Base and CodeT5 are more cost-effective and can be deployed locally. Additionally, since CodeT5 is specifically pre-trained for programming languages, and PDFObj IR represents program-like structures, CodeT5 outperforms the general BERT model in PDF malware classification tasks. In the pre-training mode of PDFObj2Vec, the BERT scheme, which is specifically pre-trained for PDFObj IR, outperforms both the general BERT Base and CodeT5 models in downstream tasks. Although it slightly lags behind text-embedding-3 in overall accuracy, it performs better against adversarial attacks, a point we will discuss in the next subsection. Furthermore, we also explored the performance of directly embedding the raw content of PDF objects using these general embedding models. We evaluated their classification performance under the same experimental setup, and we will discuss these results in our ablation study (see $\S$6.6).
@@ 6.5. Adversarial Attack
Gradient-Based Attack We adapted gradient-based white-box attack method (GradArgmax) from graph classification tasks (dai2018adversarial; wan2111adversarial) for PDF classification based on AORG. GradArgmax attack uses classifier gradient information to generate adversarial samples, aiming to reduce the model’s confidence in classifying the target sample. This approach evaluates the robustness of the integrated PDFObj2Vec graph classifier under fine-tuned adversarial examples. In this scenario, we assume the attacker has full access to the model’s gradient information, with no perturbation distance constraints. The attack budget is set to 1000, consistent with other graph attack study (wan2111adversarial). The goal is to identify which graph edges most influence the target classification by evaluating their effect on the loss function using greedy strategy. Specifically, for each candidate edge, its gradient value is computed: if negative, the edge is considered for removal; if positive, it is considered for addition. Since the number of nodes and edges in AORG varies, smaller graphs may require only slight perturbations, while larger graphs need more significant ones. To address this, we introduce the relative perturbation ratio (RPR), defined as the ratio of perturbations to the maximum number of edges in the AORG.
@@
Genetic Algorithm-Based Attack We implemented a genetic algorithm-based black-box attack (GeneticAlg) (dai2018adversarial), where the attacker has no access to the model’s internal details, only the output confidence scores. GeneticAlg leverages a genetic algorithm to optimize the structure of the input graph, effectively bringing the sample closer to the model’s decision boundary. This approach is used to evaluate the defensive capabilities of our proposed graph classifier against adversarial samples generated through complex search strategies. And, GeneticAlg consists of five key components: population, fitness function, selection, crossover, and mutation. The attack involves edge flipping, node injection, and deletion. Each generation evolves based on the fitness of the mutated samples, guiding the direction for subsequent generations. Following the experimental setup in (dai2018adversarial), we set the attacker query limit to 1000, the population size to 100, and the number of generations to 10. The results, shown in Figure9(b), indicate that Word2Vec-20k and PV-DM-20k performed the worst in terms of TRA, while the other models maintained TRA values above 0.98, with text-embedding-3 and BERT-65k performing the best. Overall, the results suggest that our proposed classifier demonstrates strong defenses against the GeneticAlg attack.
@@
Random Noise Attack We extended node classification attack methods from (ma2020towards) to perturb node features for whole-graph classification. Specifically, we injected Gaussian noise into the node features of the AORG graph to evaluate the model’s resistance to random perturbations, thereby testing its robustness under noisy conditions. To select the attack nodes, we employed degree centrality, as it provides a reasonable baseline for graph-based attacks (newman2018networks). High-degree nodes were prioritized because they play a crucial role in information propagation and are more likely to significantly impact the overall graph structure. Gaussian noise was then introduced to perturb the features of these nodes. In this black-box attack scenario, the attacker only has access to the model’s binary output (0 or 1) and is limited to a maximum of 1000 queries. The experimental results, shown in Figure 9(c), indicate that Word2Vec exhibited the lowest robustness under these conditions. In contrast, BERT-65k and BERT-20k showed the highest performance, with TRA values of 1 and 0.999, respectively. It suggests that BERT scheme are more resilient to adversarial noise, particularly with respect to node feature perturbations. The TRA of text-embedding-3 showed a significant decline under the random noise attack, compared to the GradArgmax and GeneticAlg attacks. This indicates that its embeddings are highly susceptible to noise interference.
@@
Reverse Mimicry Attack We used the reverse mimicry adversarial attack (maiorca2013looking) for evaluation, which has been evaluated in previous state-of-the-art robust PDF malware classifiers (chen2020training; tong2019improving). This black-box attack is independent of specific classifiers or features, making it suitable for evaluating our proposed graph classifier as well as the RTC and CFTC classifiers, which employ different models and features. Unlike other attacks, reverse mimicry directly manipulates the sample space by injecting malicious payloads into benign samples, thereby generating realizable adversarial examples designed to evade classification boundaries. This approach enables a rigorous evaluation of the model’s robustness against intentionally disguised adversarial samples.
@@
We followed the adversarial evaluation settings from (chen2020training; tong2019improving), using 500 seed samples to create adversarial samples. A Cuckoo sandbox (Cuckoo) was set up to test the adversarial examples. If they generated the expected network communication metrics, they were deemed legitimate. The experimental results, shown in Figure 10, reveal that with a 20k corpus size, only BERT-20k demonstrated some resistance to adversarial samples. However, with a 65k corpus size, all three PDFObj2Vec pre-training models successfully detected adversarial samples. BERT-65k achieved a TRA of 1, while the other models had TRA values below 0.2. These findings suggest that BERT-65k is particularly effective at capturing the semantic information of nodes. When malicious payloads are injected into benign samples, the corresponding malicious nodes are also present in the adversarial samples. BERT-65k accurately represents these embeddings and, through the node propagation and aggregation process in GIN, successfully identifies the adversarial samples. Notably, RTC and CFTC achieved a TRA of over 0.5 but at the cost of significantly reduced classification accuracy, as shown in Table 4. Specifically, CFTC achieved only 97.58% accuracy on the baseline dataset. And both CFTC and RTC exhibited significantly lower TNR compared to our proposed classifier. This indicates a higher FPR for CFTC and RTC, for instance, CFTC reached an FPR of 4.96% on the baseline dataset and an FPR of 12.59% on the extended dataset, substantially higher than that of BERT-65k.
@@
Results Analysis We evaluated our proposed classifier using four adversarial attack methods that target different space. Among these, the reverse mimicry attack operates in the sample space and simulates a realistic black-box scenario. In contrast, GradArgmax, GeneticAlg, and RandomNoise are feature-space attacks that, while less realistic, are widely used to stress-test classifiers under strong assumptions (dai2018adversarial; wan2111adversarial; ma2020towards; chen2020training). These attacks help reveal how vulnerable a model may be when its decision boundaries are exploited, offering valuable insights into worst-case robustness.